OpenStudy (anonymous):

Ignore air resistance. Consider a knuckleball with lateral motion satisfying the initial value problem x’’(t) = -25 sin (4ωt + θ0), x’(0) = x(0) = 0. Whit θ0 = 0 and ω = 1, find an equation for x(t) and graph the solution for 0 ≤ t ≤ 0.68. (HINT: Use integration by subtitution)

OpenStudy (anonymous):

\[x''(t)=-25\sin(4\omega t+\theta_0)\] With the initial conditions \(x'(0)=x(0)=0\). Also, you're given actual values of \(\omega\) and \(\theta_0\), so you have the differential equation \[x''(t)=-25\sin(4t)\] Integrate both side with respect to \(t\): \[\int x''(t)~dt=\int-25\sin(4t)~dt\\ x'(t)=-25\left(\cdots\right)+C_1\] Integrate again: \[\int x'(t)~dt=\int\left[-25\left(\cdots\right)+C_1\right]~dt\\ x(t)=\cdots\] I'll let you fill in the blanks. Your general solution will have some constants, \(C_1\) and \(C_2\). You solve for these using the initial values.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would that be 16 in the blanks?

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

\[\begin{align*}x'(t)&=\int(-25\sin(4t))~dt\\ &=-25\int\sin(4t)~dt\\ &=-25\left(-\frac{1}{4}\cos(4t)+C_1\right)\\ &=\frac{25}{4}\cos(4t)+C_1 \end{align*}\] Integrate both sides again, and you have \(x(t)=\cdots\). Now, use \(x'(0)=0\): \[x'(0)=0=\frac{25}{4}\cos(4\cdot0)+C_1~\Rightarrow~C_1=-\frac{25}{4}\] Do the same with \(x(0)=0\) to solve for \(C_2\) in \(x(t)=\cdots\).

OpenStudy (anonymous):

@rizwan_uet can u help?

OpenStudy (anonymous):

i already did what sith told me and i got x(t) = 25/16 sin(4t) for the first part..but i dont get when he said do the same with x(0) = 0 to solve for c2 in (t)

OpenStudy (anonymous):

ok let me check

OpenStudy (anonymous):

so what you got when you integrated it again???

OpenStudy (anonymous):

i got 25/16 sin(4t) + c1

OpenStudy (anonymous):

isnt it become x(t)= 25/16 sin(4t) + c2 ????? you put c1 here but its c2 actualy

OpenStudy (anonymous):

ohhh ok

OpenStudy (anonymous):

puting x(0) = 0 in this equation now you can get c2 thus 0 = 25/16sin(4x0) +c2

OpenStudy (anonymous):

ohhh okay i see.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are welcome